# Simplifying Hypergraph Neural Networks

### v2

@@ Abstract
Hypergraphs, with hyperedges connecting multiple nodes, are crucial for modelling higher-order interactions in real-world data. In frameworks utilising hypergraphs for downstream tasks, a task-specific model is typically paired with a hypergraph neural network (HNN). HNNs enhance the task-specific model by generating node features with hypergraph structural information via message passing. However, the training for HNNs is often computationally intensive, which limits their practical use. To tackle this challenge, we propose an alternative approach by integrating hypergraph structural information into node features using a training-free model called simplified hypergraph neural network (SHNN) that only contains a predefined propagation step. We theoretically show the efficiency and effectiveness of SHNN by showing that: 1) It largely reduces the training complexity when solving hypergraph-related downstream tasks compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions. Experiments in node classification and hyperedge prediction showcase that, compared to state-of-the-art HNNs, SHNN leads to both competitive performance and superior training efficiency. Notably, on Cora-CA, the SHNN-based framework achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
@@ 1 Introduction
Higher-order interactions involving more than two entities exist in various domains, such as co-authorships in social science (Han et al., 2009) and the spreading phenomena in epidemiology (Jhun, 2021). To model these complicated interactions, hypergraphs, an extension of traditional graphs, are widely employed (Bick et al., 2023). Hypergraphs consist of nodes representing entities and hyperedges denoting higher-order interactions that connect multiple entities. Recently, hypergraph neural networks (HNNs) have leveraged message passing over hypergraph structures to enhance node feature learning, achieving remarkable performance in various downstream tasks like node classification (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023). However, the training of HNNs is typically computationally intensive, which limits their practical applications, especially in semi-supervised learning that utilises both training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
To address this challenge, we take inspiration from the literature on scalable and efficient graph neural networks (GNNs). In this literature, SIGN (Frasca et al., 2020) proposes to move the integration of graph information from the training stage to the data preprocessing phase by a training-free model, and SGC (Wu et al., 2019) builds an efficient architecture by collapsing multiple layers of graph convolutions with the intermediate non-linear activation functions removed. We aim to explore similar ideas for hypergraphs.
@@
In this paper, we propose a novel model called simplified hypergraph neural network (SHNN), by simplifying four representative HNNs in two steps. First, we make the selected HNNs training-free by eliminating their learnable parameters. Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer HNN into a single propagation step. Remarkably, this two-step simplification unifies the four chosen HNNs into a single formulation, despite their different underlying design philosophies. We refer to this formulation as SHNN. We further conduct a thorough analysis of SHNN to theoretically demonstrate its efficiency and effectiveness. Generally, we present three key properties of SHNN: 1) SHNN significantly reduces training complexity in hypergraph-related downstream tasks compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of an SHNN and an HNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range informations.
@@
$\bullet$ We develop SHNN, an efficient and effective training-free model for generating node features with the structural information in pre-defined hypergraphs.
$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range interactions.
$\bullet$ We conduct extensive experiments in both node classification and hyperedge prediciton tasks to compare SHNN with nine state-of-the-art HNNs. The empirical results show that the SHNN-based framework exhibits both the competitive performance and superior training efficiency.
Figure 1: Training pipeline of an HNN vs. an SHNN in node classification. Top row: The HNN uses a hypergraph structure to generate node features by $(L+1)$ learnable forward propagation blocks (FP Blocks), which are then used by a classifier, with the HNN and the classifier being trained simultaneously. Bottom row: The SHNN-based framework consists solely of the classifier being trained for node classification and the SHNN, which can be precomputed prior to classifier training.
@@ 2 Notation
Hypergraphs. Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ be a hypergraph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ is the hyperedge set, and $\mathbf{H}\in\{0,1\}^{n\times m}$ is an incidence matrix in which $\mathbf{H}_{ik}=1$ means that $e_{j}$ contains node $v_{i}$ and $\mathbf{H}_{ik}=0$ otherwise. Define $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal matrix of node degrees and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}\in\mathbb{R}_{\geq 0}^{m\times m}$ as a diagonal matrix of hyperedge degrees, where $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}_{ii}}$ and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}$ are the number of hyperedges with $v_{i}$ and the number of nodes in $e_{k}$, respectively. In the following discussions, we assume the hypergraph does not have isolated nodes or empty/duplicate hyperedges. Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less; and 3) The neighbouring nodes of a node are its 1-hop neighbours.
Graphs. Let $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ be a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, and $\mathbf{W}\in\mathbb{R}_{\geq 0}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$ in which $\mathbf{W}_{ij}>0$ means that $v_{i}$ and $v_{j}$ are connected and $\mathbf{W}_{ij}=0$ otherwise. We set $\mathbf{D}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal node degree matrix for $\mathcal{G}$, where $\mathbf{D}_{ii}$ is the sum of the $i$-th row of $\mathbf{W}$. Moreover, we denote the graph Laplacian of $\mathcal{G}$ as $\mathbf{L}=\mathbf{D}-\mathbf{W}$.
Clique expansion. Given a hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, its clique expansion is defined as a graph $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$, where $\mathcal{V}$ remains unchanged, and $\mathbf{W}_{ij}>0$ if and only if $v_{i}$ and $v_{j}$ are connected by a hyperedge on $\mathcal{H}$ and $\mathbf{W}_{ij}=0$ otherwise. Hence, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
@@ 3 Methodology
In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models along with their corresponding learning frameworks for downstream tasks into general formulations. Furthermore, we develop the simplified hypergraph neural network (SHNN) by removing the learnable parameters and the non-linear function from four representative HNNs.
3.1 Revisiting Hypergraph Neural Networks
@@
Table 1: Overview for four state-of-the-art HNNs. In this table, $\gamma_{U},\gamma_{E},\gamma_{D},\gamma^{\prime}_{l}\!\in\!(0,1)$ are hyperparameters and $\mathbf{I}\in\mathbb{R}^{d\times d}$ denotes an identity matrix.
Name Type Message Passing Formulation
@@
UniGCNII (Huang & Yang, 2021) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$
@@
Deep-HGNN (Chen et al., 2022) Direct Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$
AllDeepSets (Chien et al., 2022) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$
ED-HNN (Wang et al., 2023a) Indirect Message Passing $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$
@@
Model design. Generally, existing HNNs are based on the message passing mechanism, adapted from the graph machine learning literature (Xu et al., 2019; Bronstein et al., 2021). This mechanism makes models iteratively update the features of each node by aggregating the features of its neighbours. Mathematically, the features of a specific node $v_{i}$ are updated by the following propagation operators:
@@
where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable non-linear function that facilitates the message passing between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of $1$-hop neighbours of $v_{i}$ on the hypergraph. Specifically, $\phi_{\mathbf{\Theta}}(\cdot)$ can be categorized into two types: direct and indirect message passing functions. The direct approaches aggregate features of neighbouring nodes to that of the target node directly (Huang & Yang, 2021; Wang et al., 2023b; Chen et al., 2022). In the indirect approaches, the features of neighbouring nodes are aggregated to that of the target node via some virtual nodes (Chien et al., 2022; Wang et al., 2023a; Yang et al., 2022). Mathematically these approaches can be formulated as:
$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{G}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (2a)
@@
where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable non-linear functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a learnable vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$. Here, $n_{v}$ is a hyperparameter corresponding to the number of the virtual nodes, and Eq. (2a) and Eq. (2b) represent the direct and indirect approaches respectively. Further, we detail four representative HNNs in Table 1, and show how specific HNNs are special cases of Eq. (1) in Appendix.
Application for downstream tasks. The pipeline of an HNN-based learning framework for downstream tasks can be formulated as follows:
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific model, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^1] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by an HNN, and $\Phi_{\mathbf{\Theta}}(\cdot)$ is an HNN that updates the node features with the forward propagation functions formulated as Eq. (1). The models $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks. As illustrated in Table 1, existing HNNs contain complex learnable message passing formulation, which makes the training of the HNN-based framework computationally expensive.
3.2 Simplifying Hypergraph Neural Networks
In this section, our goal is to introduce the simplified hypergraph neural network (SHNN) by simplifying the representative HNNs presented in Table 1.
Learnable parameters elimination. To develop a training-free model, we first eliminate the learnable parameters from the message passing formulation of four selected HNNs. Specifically, we replace the learnable matrices in their formulations with identity matrices. This simplification allows us to represent the simplified formulations of the chosen HNNs as follows:

### v3

@@ Abstract
Hypergraphs are crucial for modeling higher-order interactions in real-world data. Hypergraph neural networks (HNNs) effectively utilise these structures by message passing to generate informative node features for various downstream tasks like node classification. However, the message passing block in existing HNNs typically requires a computationally intensive training process, which limits their practical use. To tackle this challenge, we propose an alternative approach by decoupling the usage of the hypergraph structural information from the model training stage. The proposed model, simplified hypergraph neural network (SHNN), contains a training-free message-passing block that can be precomputed before the training of SHNN, thereby reducing the computational burden. We theoretically support the efficiency and effectiveness of SHNN by showing that: 1) It is more training-efficient compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions. Experiments based on six real-world hypergraph benchmarks in node classification and hyperlink prediction present that, compared to state-of-the-art HNNs, SHNN shows both competitive performance and superior training efficiency. Specifically, on Cora-CA, SHNN achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
@@ 1 Introduction
Higher-order interactions involving more than two entities exist in various domains, such as co-authorships in social science (Han et al., 2009) and the spreading phenomena in epidemiology (Jhun, 2021). Hypergraphs, extending traditional graphs, effectively model these complex interactions with nodes representing entities and hyperedges connecting multiple entities (Bick et al., 2023). To leverage such structures for solving downstream tasks, the design of learning algorithms on hypergraphs has garnered increasing attention.
Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most current research focuses on developing hypergraph neural networks (HNNs) with a message passing block (MP-Block) and a task-specific block [^1] For instance, in node classification, the task-specific block is implemented as a classifier, whereas in hyperlink prediction, it is designed as a predictor. Following previous works (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023), we focused on designing MP-Blocks that are compatible with various task-specific blocks.. The MP-Block enables information exchange between connected nodes to generate informative node features for the task-specific block (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023). However, similar to other message passing neural networks (Frasca et al., 2020; wu2022graph), training the MP-Block makes loss computation interdependent for connected nodes, resulting in a computationally intensive training process for existing HNNs. This limits their practical applications, especially in semi-supervised learning that utilises the training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
@@
To address this challenge, inspired by the literature on efficient and effective GNNs (Frasca et al., 2020; Wu et al., 2019), our key solution is to move the integration of the hypergraph structural information from the training stage to the data preprocessing phase. This leads to a novel model called simplified hypergraph neural network (SHNN). In this model, a task-specific block is paired with a simplified message passing block (SMP-Block) that is training-free and can be precomputed in the data preprocessing phase. We develop the SMP-Block by simplifying the MP-Block of four state-of-the-art HNNs (Huang & Yang, 2021; Chen et al., 2022; Chien et al., 2022; Wang et al., 2023a) in two steps. First, we make the selected MP-Blocks training-free by eliminating their learnable parameters. Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer MP-Blocks into a single propagation step. Remarkably, this two-step simplification unifies the chosen MP-Blocks into a single formulation, despite their different underlying design philosophies. We refer to this formulation as SMP-Block. Further, we thoroughly analyse SHNN to demonstrate its efficiency and effectiveness theoretically. Generally, we show three key properties: 1) SHNN is more training-efficient compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of the HNN and our SHNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range information.
@@
$\bullet$ We develop SHNN, an efficient and effective model for hypergraph-structured data. To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, significantly enhancing training efficiency.
$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range information.
$\bullet$ We conduct extensive experiments in both node classification and hyperedge prediciton tasks to compare SHNN with nine state-of-the-art HNNs. The empirical results show that the proposed SHNN exhibits both the competitive performance and superior training efficiency.
Figure 1: Training pipeline of HNN vs. SHNN in node classification. Top row: HNN uses a hypergraph structure to generate node features by a learnable MP-Block, which are then used by a classifier, with the MP-Block and the classifier being trained together. For brevity, we omit the MLP in HNN for the input node features. Bottom row: SHNN consists solely of the classifier being trained for node classification and the SMP-Block, which can be precomputed prior to classifier training.
@@ 2 Notation
Hypergraph. Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ be a hypergraph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ is the hyperedge set, and $\mathbf{H}\in\{0,1\}^{n\times m}$ is an incidence matrix in which $\mathbf{H}_{ik}=1$ means that $e_{j}$ contains node $v_{i}$ and $\mathbf{H}_{ik}=0$ otherwise. Define $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal matrix of node degrees and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}\in\mathbb{R}_{\geq 0}^{m\times m}$ as a diagonal matrix of hyperedge degrees, where $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}_{ii}}$ and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}$ are the number of hyperedges with $v_{i}$ and the number of nodes in $e_{k}$, respectively. In the following discussions, we assume the hypergraph does not have isolated nodes or empty/duplicate hyperedges. Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; and 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less.
Graph and clique expansion. Let $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ be a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, and $\mathbf{W}\in\mathbb{R}_{\geq 0}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$ in which $\mathbf{W}_{ij}>0$ means that $v_{i}$ and $v_{j}$ are connected and $\mathbf{W}_{ij}=0$ otherwise. We set $\mathbf{D}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal node degree matrix for $\mathcal{G}$, where $\mathbf{D}_{ii}$ is the sum of the $i$-th row of $\mathbf{W}$. Moreover, we denote the graph Laplacian of $\mathcal{G}$ as $\mathbf{L}=\mathbf{D}-\mathbf{W}$. Given a hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, its clique expansion is defined as a graph $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$, where $\mathcal{V}$ remains unchanged, and $\mathbf{W}_{ij}>0$ if and only if $v_{i}$ and $v_{j}$ are connected by a hyperedge on $\mathcal{H}$ and $\mathbf{W}_{ij}=0$ otherwise. Therefore, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
@@ 3 Methodology
In this section, we aim to develop an efficient and effective model for hypergraph-structured data. In Subsection 3.1, we present the overall design of our simplified hypergraph neural network. Subsection 3.2 elaborates on the development of the simplified message passing block.
3.1 Overview
General formulation. Inspired by a previous work named SIGN (Frasca et al., 2020) in the GNN literature, we design a model, named simplified hypergraph neural network (SHNN), that can integrate the information about hypergraph structures into the data preprocessing stage, thus also removing the reliance on it at training. Mathematically, the SHNN can be summarised as the following formulation:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=S(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (1)
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^2] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $S(\cdot)$ denotes an MP-Block that contains only pre-defined parameters, allowing it to be precomputed before the training phase, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $S(\cdot)$. In SIGN, $S(\cdot)$ is implemented by the combination of some graph filters. However, we cannot directly use existing hypergraph filters (Zhang et al., 2019; Qu et al., 2022) for two main reasons. Firstly, existing hypergraph filters are primarily designed for $k$-uniform hypergraphs, where all hyperedges have size $k$. Secondly, these filters require the use of an incidence tensor, which leads to prohibitively high memory usage for practical applications. Specifically, for a hypergraph with $n$ nodes and hyperedges of uniform size $k$, the corresponding incidence tensor has $n^{k}$ dimensions. Hence, instead of using existing methods, we draw inspiration from SGC (Wu et al., 2019) in the GNN literature to design a novel $S(\cdot)$ for hypergraphs, named simplified message passing block (SMP-Block), by simplifying MP-Blocks in state-of-the-art HNNs.
3.2 Simplified Message Passing Block
@@
In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models and their MP-Blocks into general formulations. Furthermore, we develop the simplified message passing block (SMP-Block) by removing the learnable parameters and the non-linear function from the MP-Block in four state-of-the-art HNNs.
3.2.1 Revisiting Hypergraph Neural Networks
@@
Table 1: Overview for four state-of-the-art HNNs and our SHNN. In this table, $\gamma_{U},\gamma_{E},\gamma_{D},\gamma^{\prime}_{l}\!\in\!(0,1)$ are hyperparameters and $\mathbf{I}\in\mathbb{R}^{d\times d}$ denotes an identity matrix. Moreover, $M$ is the training computational complexity of the task-specific block, $n$ is the node count, $m$ is the hyperedge count, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero values in $\mathbf{H}$, $T$ is the number of training epochs, $L$ is the number of layers, and $d$ is the feature dimension.
Name Type Hypergraph-Wise MP Function Training Computational Complexity
@@
UniGCNII (Huang & Yang, 2021) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$
@@
Deep-HGNN (Chen et al., 2022) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$
AllDeepSets (Chien et al., 2022) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
ED-HNN (Wang et al., 2023a) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
SHNN (Ours) Direct MP $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$ $\mathcal{O}(M)$
General formulation. An HNN for downstream tasks is formulated as the following equation:
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent, $\Phi_{\mathbf{\Theta}}(\cdot)$ is a learnable message passing block, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $\Phi_{\mathbf{\Theta}}(\cdot)$. The blocks $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks. We detail the message passing block in the next paragraph.
@@
Message passing block. Generally, existing HNNs are based on the message passing mechanism, adapted from the graph machine learning literature (Xu et al., 2019; Bronstein et al., 2021). This mechanism makes models iteratively update the features of each node by aggregating the features of its neighbours. Mathematically, the features of a specific node $v_{i}$ are updated by the following propagation operators:
@@
where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable node-wise message passing function that facilitates the information exchange between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of neighbours of $v_{i}$ on the hypergraph. Specifically, $\phi_{\mathbf{\Theta}}(\cdot)$ can be categorized into two types: direct and indirect message passing functions. The direct approaches aggregate features of neighbouring nodes to that of the target node directly (Huang & Yang, 2021; Wang et al., 2023b; Chen et al., 2022). In the indirect approaches, the features of neighbouring nodes are aggregated to that of the target node via some virtual nodes (Chien et al., 2022; Wang et al., 2023a; Yang et al., 2022). Mathematically these approaches can be formulated as:
$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (4a)
@@
where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$. Here, $n_{v}$ is a hyperparameter corresponding to the number of the virtual nodes, and Eq. (4a) and Eq. (4b) represent the direct and indirect approaches respectively. Further, we summarise four state-of-the-art HNNs in Table 1, and provide a more detailed review of existing HNNs in Appendix. In HNNs, an MP-Block typically consists of a combination of $L$ message-passing functions. In the following subsection, we introduce our SMP-Block by simplifying the MP-Blocks of HNNs in Table 1.
3.2.2 Simplifying Hypergraph Neural Networks
Our goal is to introduce the SMP-Block by simplifying the state-of-the-art HNNs presented in Table 1.
Learnable parameters elimination. To develop a training-free MP-Block, we first eliminate the learnable parameters from the message passing functions of the selected HNNs. Specifically, we replace the learnable matrices in these functions with identity matrices. This simplification allows us to represent the simplified message passing functions as follows: